Two massless strings connect a sphere of mass 9.76kg to a vertical pole where d=1.17m and l=2.33m.The spheres revolves in a horizontal circle at a constant speed of 9.33 m/s (without the wires wrapping around the pole). Determine the magnitudes of the tension in each string.




1
Expert's answer
2021-11-23T10:33:08-0500


The centripetal force is balanced by the tensions:


"F_c=m\frac{v^2}r=T_1+T_2,\\\space\\
T_1=m\frac{v^2}r=726\text{ N},\\
T_2=0."